OpenStudy (anonymous):

please help!! Which choice could be the equation of a line perpendicular to the line represented by this equation? Be sure to explain your solution. y = (2/5)x + 2 a. y = (2/5)x - 3 b. y = (5/2)x + 7 c. y = (-5/2)x - 3 d. y = -2x + 5

jimthompson5910 (jim_thompson5910):

what is the slope of the given line

OpenStudy (anonymous):

im not really sure how to work this

jimthompson5910 (jim_thompson5910):

what is the given line first off

jimthompson5910 (jim_thompson5910):

what is the equation of the given line

OpenStudy (anonymous):

y = (2/5)x + 2?

jimthompson5910 (jim_thompson5910):

good, would you agree that it's in slope intercept form?

OpenStudy (anonymous):

yes it has to be

jimthompson5910 (jim_thompson5910):

what is slope intercept form in general

OpenStudy (anonymous):

(2/5)x?

jimthompson5910 (jim_thompson5910):

is slope intercept form Ax+By = C? is it y = mx + b? or is it y-y1 = m(x-x1)

OpenStudy (anonymous):

y=mx+b

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what does m represent?

OpenStudy (anonymous):

slope

jimthompson5910 (jim_thompson5910):

so the slope is the number or fraction in front of the x

OpenStudy (anonymous):

yees

jimthompson5910 (jim_thompson5910):

in the case of y = (2/5)x + 2, the slope is 2/5 the x is NOT part of the slope

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

to find the slope of any perpendicular line to this one, we flip the fraction and flip the sign flip the fraction: 2/5 ----> 5/2 flip the sign: 5/2 ----> -5/2

jimthompson5910 (jim_thompson5910):

so if the original slope is 2/5, then the perpendicular slope is -5/2

jimthompson5910 (jim_thompson5910):

so which equation has a slope of -5/2 (in your answer choices)?

OpenStudy (anonymous):

C

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

that's your answer

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome